OpenStudy (anonymous):

~ simplify the following expressions : a)5(x+3)+3(2x-3) = ? b)3(x-4)+2(4x-1) = ? c)2(5x-3)+3(4-2x) = ? d)3(4x-3)+5(x+2) = ? e)2(4x+3)-3(x+2) = ? f)6(2x-1)-3(x-3) = ?

OpenStudy (vishweshshrimali5):

a) 2x + 15 + 6x - 9 = 8x + 6 b)3x - 12 + 8x - 2 = 11x - 14 c) 10x -6 +12 - 6x = 4x + 6 d) 3(4x-3)+5(x+2) = 12x - 9 + 10x + 10 = 22x + 1 e) 2(4x+3)-3(x+2) = 8x + 6 - 3x - 6 = 5x f) 6(2x-1)-3(x-3) = 12x - 6 - 3x + 9 = 9x + 6

OpenStudy (vishweshshrimali5):

i hope I am correct

OpenStudy (anonymous):

check the first once more

OpenStudy (diyadiya):

1. 11x+6

OpenStudy (anonymous):

a) what is it ?

OpenStudy (anonymous):

??

OpenStudy (diyadiya):

a) 11x+6